OpenStudy (anonymous):

Solve by factoring 6t^2+t-5=0

OpenStudy (amtran_bus):

t(6t+1)-5=0

OpenStudy (freckles):

6(-5)=-30 and 6+(-5)=-1 so 6t^2+6t-5t-5=0 factor the left hand side by grouping

OpenStudy (anonymous):

Thank you ! I get it now